1) Two hundred randomly selected women in a certain city, including those who are working, were asked the time they spend in the beauty salon for each visit. The average time was 44 minutes with a standard deviation of 16 minutes. Construct an 80% confidence interval for the mean time spent in the salon by all the women in the city.

a.[43.05, 44.95]

b.[43.76, 44.24]

c.[42.55, 45.45]

d.[43.64, 44.36]

e.[42.45, 45.55]

2) In order to estimate the mean number of books purchased by customers, a bookstore samples 64 customers and obtains a sample mean of 6 books with a sample standard deviation of 2. Construct a 95% confidence interval for the mean number of books purchased by all the customers visiting the store.

a.[5.51, 6.49]

b.[5.71, 6.29]

c.[5.59, 6.41]

d.[5.65, 6.35]

e.[5.50, 6.50]
Answer #1

Solution : ( 1 )

c. (42.55, 45.45)

Explanation

80% confidence interval for the mean time spent in the salon by all the women in the city

-(4-12,. 720 12.20 44-1.28 * 44+1.28 * V200

44 1.28 V23 * 52

=( 20, 22 ) 44-1.28 * 44 + 1.28 *

=left ( 44-1.28* rac{2^4}{2^{rac{3}{2}}*5},quad:44+1.28* rac{2^4}{2^{rac{3}{2}}*5} ight )

44-1.28 * 44 + 1.28 *-

=left ( 44-1.28* rac{2^2sqrt{2}}{5},quad:44+1.28* rac{2^2sqrt{2}}{5} ight )

4V2 44 1.28 * 44+1.28

=left ( 44- rac{1.28*4sqrt{2}}{5},quad:44+ rac{1.28*4sqrt{2}}{5} ight )

5.12V2 5.12V2

-(«-32«214) 5.12 1.414 5.12 1.414

=left ( 44- rac{7.24077}{5},quad:44+ rac{7.24077}{5} ight )

=left ( 44- 1.448154,quad:44+ 1.448154 ight )

42.551846, 45.448154)

(42.55, 45.45)​​​​​​​

======================================================================================

Solution : ( 2 )

mathrm{a.quad:}left ( 5.51,quad:6.49 ight )

Explanation

mathrm{95%:confidence:interval:for:the:mean:number:of:books:purchased:by:all:the:customers:visiting:the:store::}

(6-196. *石, 6+1.96 * 6+196 V64

=left ( 6-1.96* rac{2}{8},quad:6+1.96* rac{2}{8} ight )

6-1.96 *-, 6+1.96 *-

=left ( 6-rac{1.96}{4},quad:6+rac{1.96}{4} ight )

=left ( 6-0.49,quad:6+0.49 ight )

=left ( 5.51,quad:6.49 ight )